# Enhancing High-Level Synthesis with Automated Pragma Insertion and Code Transformation Framework

### v3

@@ 6.3.1. AutoDSE and NLP-DSE
bert_100_3072 1 D,P 220.29 1.10x 1.10x
bert_100_3072 0 D,P 219.97 1.11x 1.11x
gemm 1 D,P 198.26 1.79x 1.51x
gemm 0 D,P 210.63 9.25x 9.57x
bicg 1 D,P 1.96 1.97x 1.97x
bicg 0 D,P 1.86 1.92x 1.93x
gemver 1 P 17.85 4.64x 1.77x
gemver 0 P 12.45 30.91x 7.24x
mvt 1 P 13.71 1.76x 1.76x
mvt 0 P 10.00 1.45x 1.45x
bert_3072_100 1 D,P,T 42.93 2.98x 3.31x
bert_3072_100 0 D,P,T 44.67 536.42x 626.60x
cnn 1 D,P,T 340.20 8.00x 8.91x
cnn 0 D,P,T 339.34 8.05x 10.67x
doitgen 1 D,P 53.22 1.33x 2.66x
doitgen 0 D,P 54.05 5.46x 2.71x
symm 1 D,P 236.82 8.15x 7.20x
symm 0 D,P 236.56 8.14x 7.19x
syr2k 1 D,P 428.12 9.28x 3.12x
syr2k 0 D,P 428.39 17.57x 3.13x
syrk 1 D,P 306.07 12.44x 4.40x
syrk 0 D,P 308.71 12.55x 4.43x
Average 153.65 59.99x 22.85x
Geo Mean 69.09 5.97x 2.89x
Table 3. Comparison of the Throughput (GF/s) achieved with Sisyphus, AutoDSE and NLP-DSE
Sisyphus averages a speedup of 59.99x and 22.85x over AutoDSE and NLP-DSE, respectively, across the evaluated kernels. In terms of geometric mean, Sisyphus achieves a speedup of 5.97x and 2.89x.
Conserve the original schedule when needed  In the case of atax and jacobi-2d, Sisyphus preserves the original schedule and solely inserts pragmas. Thus, it effectively maintains the existing schedule while efficiently incorporating pragmas, a process comparable to that of two DSEs dedicated to pragma insertion. Regarding heat-3d, the original design opts for partial unrolling of the three loops with unroll factors of 2, 2, and 38, respectively, enabling a level of parallelism unattainable without code transformation. However, upon evaluation, we found that this design was over-utilizing resources. Therefore, the NLP solution, using a pessimistic constraint for DSP reuse, resulted in a design with the original schedule.
Although we note improvements for jacobi-2d, it is noteworthy that the same pragma insertion was utilized in at least one DSE. However, despite this consistency, AMD/Xilinx Vitis 2023.2 failed to synthesize the design generated by the source-to-source compiler Merlin, which the two DSEs employ.
Code transformation to manage reduction loop  In the cases of 2mm, bert, and gemm, we observe notable improvements resulting from code transformations, particularly in effectively managing the reduction loop. This enhancement is particularly evident with bert_3072_100, where the reduction loop exhibits a large trip count, as well as in gemm when tree-reduction is not used.
Code transformation with tiling  The selection of tile size becomes crucial for determining which array sizes, particularly for bert_3072_100 and CNN with their large arrays, should be cached on-chip and where they should be transferred. Organizing the outermost loop as output stationary (cf. Section 5.3), prompts the NLP to perform partial transfers of the array iterated by these loops in these examples. Consequently, the NLP’s choices in these scenarios aim to optimize data reuse and minimize redundant memory transfers. However, for CNN, its large size requires multiple transfers of the array input (cf. Listing 4).
Code transformation for memory-bound kernel  For mvt and gemver, both memory-bound kernels, effective optimization entails loop permutations to prevent redundant array transfers and accommodate array partitioning within the constraints of the AMD/Xilinx compiler’s 1024 limit. Sisyphus adeptly manages these loop permutations and pragma insertions, guaranteeing optimized performance for these kernels. Additionally, in the case of gemver without tree reduction, the NLP reverses all loop orders to pipeline the non-reduction loop, further enhancing the QoR. For bicg, fully distributing and arranging the loops appropriately can enhance parallelization capabilities.
Allow to achieve a parallelism not achievable without loop transformation  For doitgen, symm, syrk, and syr2k, the original schedule imposes constraints on the achievable level of parallelism. This limitation elucidates the performance enhancements facilitated by Sisyphus. In the case of syrk and syr2k, the primary advantageous transformation is the maximal loop distribution. Meanwhile, for doitgen and symm, we observe QoR improvements because code transformations enable a combination of loop unrolling not attainable even with maximal loop distribution.
@@ 6.3.2. HARP
In Table 4 we compare the throughput and resource utilization (BRAM, DSP, LUT, FF) achieved with HARP (Sohrabizadeh et al., 2023). The enhanced performance seen in both bicg and gemver can be attributed to the same underlying factor observed in the comparison with AutoDSE and NLP-DSE. Regarding mvt, HARP produced comparable performance to our own due to their use of the Merlin source-to-source compiler, which facilitates loop permutations when partially unrolling consecutive loops. This allows for them to find the same schedule that our framework identified. However, for Gemm, the design discovered by HARP falls outside our specified parameter space, particularly because of our restriction on maximal array partitioning. Nonetheless, it is worth noting that it is the only design among the top 10 evaluated that demonstrates a performance enhancement over our framework.
GF/s Resource Utilization (%) Perf. Imp
@@
Kernel HARP Sisyphus HARP Sisyphus
atax 1.72 1.77 78,52,49,50 38,26,25,31 1.03x
bicg 0.92 1.77 75,25,35,36 38,26,24,26 1.92x
gemm 125.59 99.96 29,80,55,40 1,67,51,34 0.80x
gemver 1.66 7.38 29,28,35,19 43,27,39,47 4.45x
mvt 7.07 7.06 40,78,43,30 21,64,33,26 1.00x
Average 27.39 23.58 1.84x
Geo Mean 4.71 6.95 1.47x
Table 4. Comparison of GF/s and resources utilization with HARP
6.4. Latency estimation
Implementing the post-optimization strategy outlined in Section 5.2 leads to a loss of the lower bound discussed in (Pouget et al., 2024a, b), primarily because we do not incorporate computation-communication overlap modeling into the NLP. This omission is deliberate, as including such modeling would significantly increase search time.
Similarly, we opt against integrating the loop_flatten pragma into our model, despite its potential automatic invocation by the compiler for loops without reductions. As detailed in Section 5, we intentionally deactivate it for loops with reductions. Although this optimization’s automatic application results in the loss of the lower bound, we have made this choice owing to its rare implementation. Our aim is to avoid reliance on an optimization that may not be consistently available.
In Figure 1, we depict the estimated latency by the NLP of the evaluated designs and juxtapose them with the latencies reported by HLS. Our post-optimization efforts reveal that most latencies estimated by NLP exceed those reported by HLS, with the exception of CNN with and without tree reduction. In the case of CNN (the last two dots), the pragmas are implemented as anticipated, but the lower bound of the computed unrolled part is not sufficiently tight.
Figure 1. Comparison of the latency reported by HLS with the estimated latency obtained through NLP
6.5. Scalability of the NLP
Among the 42 NLP evaluations (including those with pessimistic resource reuse), only 3 timed out after 4 hours. These timeouts were specifically encountered during the bicg and atax kernels with tree reduction, and the gemver kernel without tree reduction. Throughout the search process, the NLP solver BARON maintains two bounds: a lower bound, which may be unattainable, and the lowest achievable bound discovered thus far. If the solver completes before reaching the timeout, these two bounds are identical. In the cases of atax and bicg, these two bounds were within 1% of each other, while for gemver, the difference was 20%. The average solving time was 33 minutes, while the geometric mean stood at 2.94 minutes. Out of 42 instances, 35 finished within an hour.
@@ 7. Related Work
Design Space Exploration (DSE) methodologies for pragma insertion, such as those mentioned in (Pouget et al., 2024a, b; Sohrabizadeh et al., 2021, 2023, 2022; Ferretti et al., 2018; Zhao et al., 2017; Sohrabizadeh et al., 2023; Zhong et al., 2016; Wu et al., 2023; Liu and Carloni, 2013; Mahapatra and Schafer, 2014; Yu et al., 2018), yield designs with a satisfactory QoR. However, they often require extensive computational time, with AutoDSE taking up to a day to complete. While users have the option to perform code transformations before conducting DSE, the process typically involves treating each code transformation and pragma insertion as separate optimization problems. This decomposition of the problem can result in a loss of QoR because predicting which transformation is necessary to achieve an optimal design is challenging. The NLP-DSE method, as explained in (Pouget et al., 2024b, a), employs a non-linear solver to integrate pragmas into the code. We build upon the methodology developed by the authors and tailor it to our specific optimization space to determine the pragmas and schedule for the kernel.
@@
Various code transformations have been devised for CPUs (Bondhugula et al., 2008), GPUs (Verdoolaege et al., 2013), and FPGAs (Pouchet et al., 2013; Zhao et al., 2022; Zhao and Cheng, 2021; Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018). While code transformations for CPUs and GPUs yield remarkable results tailored to their respective architectures, they may not be inherently suitable for FPGA targets aimed at achieving high parallelism. Pluto (Bondhugula et al., 2008), considered state-of-the-art, performs code transformations including tiling to minimize dependency distance between memory accesses, thereby facilitating data reuse. However, in our scenario, reducing dependency distance could potentially constrain parallelization efforts. Regarding (Zhao et al., 2022; Zhao and Cheng, 2021), they employ Pluto on different scopes of the kernel for code transformation, yet its pragma insertion capabilities are limited, making it incomparable to our work. Conversely, (Pouchet et al., 2013; Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018) pursue a distinct objective from ours. (Pouchet et al., 2013) aims to minimize communication between off-chip and on-chip, which results in better QoR than ours for memory-bound kernels. The works (Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018) concentrate on code transformations aimed at enhancing pipelining techniques. However, these objectives may not be suitable for computation-bound kernels requiring high levels of parallelization.
The (Ye et al., 2022; Lai et al., 2019; Huang et al., 2021; Chen et al., 2024; Ye et al., 2024) compilers undertake code transformation and pragma insertion. However, these transformations are somewhat limited, mostly encompassing heuristic modifications based on loop properties. Moreover, the scope of pragma insertion is more limited compared to our proposed approach.
The choice of tile sizes significantly impacts the QoR. In line with our methodology, (Liu et al., 2017a; Li et al., 2021) employ a cost model to determine the tile size. However, while (Liu et al., 2017a) focuses on minimizing communication overhead, our approach differs. On the other hand, (Li et al., 2021) investigates tiles size selection for Convolutional Neural Networks (CNNs) with three-level CPU caching.
@@ 8. Conclusion
This paper addresses challenges in optimizing HLS via code transformation and pragma insertion. While HLS and source-to-source compilers accelerate hardware design, achieving high-quality results often demands manual code intervention, prone to errors and time-consuming. To address this, Sisyphus is introduced, automating code transformation and pragma insertion for HLS into a single optimization problem. Using Non-Linear Programming (NLP), Sisyphus determines code transformations and pragma placements simultaneously, ensuring optimized performance within resource constraints.

### v4

@@ 6.3.3. AutoDSE and NLP-DSE for Bert and CNN
bert_100_3072 1 D,P 241.59 1.21x 1.21x
bert_100_3072 0 D,P 241.06 1.22x 1.22x
bert_3072_100 1 D,P,T 314.89 21.85x 24.29x
bert_3072_100 0 D,P,T 82.41 989.65x 1,156.02x
cnn 1 D,P,T 314.15 7.38x 8.23x
cnn 0 D,P,T 341.20 8.09x 10.73x
Average 212.83 103.37x 120.61x
Geo Mean 185.32 4.38x 4.69x
Table 5. Throughput (GF/s) Comparison of Sisyphus, AutoDSE and NLP-DSE Using Vitis HLS
@@ 6.3.4. HARP
In Table 6 we compare the throughput and resource utilization (BRAM, DSP, LUT, FF) achieved with HARP (Sohrabizadeh et al., 2023). The enhanced performance seen in both bicg and gemver can be attributed to the same underlying factor observed in the comparison with AutoDSE and NLP-DSE. Regarding mvt, HARP produced comparable performance to our own due to their use of the Merlin source-to-source compiler, which facilitates loop permutations when partially unrolling consecutive loops. This allows them to find the same schedule that our framework identified. However, for Gemm, the design discovered by HARP falls outside our specified parameter space, particularly because of our restriction on maximal array partitioning. Notably, it is the only design in the top 10 that outperforms our framework.
GF/s Resource Utilization (%) Perf. Impr.
@@
Kernel HARP Ours HARP Sisyphus
atax 1.72 1.96 78,52,49,50 38,65,53,33 1.14x
bicg 0.92 1.96 75,25,35,36 38,65,53,34 2.13x
gemm 125.59 88.61 29,80,55,40 32,48,24,17 0.71x
gemver 1.66 9.42 29,28,35,19 30,64,65,33 5.69x
mvt 7.07 8.24 40,78,43,30 53,43,49,30 1.17x
Average 27.39 22.04 2.16x
Geo Mean 4.71 7.66 1.62x
Table 6. Comparison with HARP Using Vitis HLS
6.4. Single SLR Onboard Evaluation
GF/s RU: BRAM,DSP,LUT(K),FF(K) Perf. Impr.
Kernel Ours Ours AutoDSE
2mm 44.32 312,549,83,94 353.5,963,284,292 126.6x
3mm 29.89 411,980,108,123 470,1117,281,306 76.6x
atax 1.03 252,169,118,79 630.5,452,170,212 4.47x
bicg 1.03 253,171,131,86 867.5,196,168,217 1.88x
Table 7. Comparison of Onboard Evaluation for 1 SLR
The resource utilization (RU) is detailed in the thousands for both LUTs and FFs. URAM utilization is excluded as no kernels use it. And the last column show the performance improvement we achieve with Sisyphus over AutoDSE. For the kernels evaluated on board, we can see that our generated design is both faster and consumes fewer resources compared to AutoDSE.
6.5. Latency Estimation and Scalability
Implementing the post-optimization strategy from Section 5 results in a loss of the lower bound discussed in (Pouget et al., 2024a, b), mainly because we exclude communication overlap modeling between loop bodies in the NLP to avoid significantly increasing search time. We also do not incorporate the loop_flatten pragma, even if the compiler may automatically apply it for loops without reductions, as we deactivate it for reduction loops. Although this leads to a lower bound loss, we prefer not to depend on an optimization that may not be consistently implemented. The average solving time of the NLP was 455 seconds, while the geometric mean stood at 6.7 seconds. Only the kernel 3mm with tree reduction timeout at 4h. 32/36 NLP’s problem are solve in less than 1 minutes. It is important to note that the solver provides a feasible solution early in the process, though it may not be theoretically optimal. As the search continues, the solution improves, and once the solver completes, the solution is theoretically optimal. Therefore, it is possible to set a timeout to obtain a feasible, but not optimal, solution more quickly.
@@ 7. Related Work
Various code transformations have been devised for CPUs (Bondhugula et al., 2008), GPUs (Verdoolaege et al., 2013), FPGAs (Pouchet et al., 2013b; Zhao et al., 2022; Zhao and Cheng, 2021; Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018) or for all (Zheng et al., 2020). While code transformations for CPUs and GPUs yield remarkable results tailored to their respective architectures, they may not be inherently suitable for FPGA targets aimed at achieving high parallelism. Regarding (Zhao et al., 2022; Zhao and Cheng, 2021), they employ Pluto on different scopes of the kernel for code transformation, yet its pragma insertion capabilities are limited, making it incomparable to our work. Conversely, (Pouchet et al., 2013b; Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018) pursue a distinct objective from ours. (Pouchet et al., 2013b) aims to minimize communication between off-chip and on-chip, which results in better QoR than ours for memory-bound kernels. The works (Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018) concentrate on code transformations aimed at enhancing pipelining techniques. However, these objectives may not be suitable for computation-bound kernels requiring high levels of parallelization. Ansor (Zheng et al., 2020) achieves impressive results for CPU and GPU architectures; however, there are notable differences with our approach. Ansor requires multiple evaluations to perform its design-space exploration (DSE) and update its cost model, which accelerates the DSE process but is not scalable for FPGA, where a single evaluation can take minutes or even hours. Furthermore, Ansor’s approach to code transformations and pragma insertions (which are more tailored for CPU/GPU) addresses these problems separately, potentially resulting in suboptimal designs. The choice of tile sizes significantly impacts the QoR. In line with our methodology, (Liu et al., 2017a; Li et al., 2021) employ a cost model to determine the tile size. However, while (Liu et al., 2017a) focuses on minimizing communication overhead, our approach differs. On the other hand, (Li et al., 2021) investigates tiles size selection for Convolutional Neural Networks (CNNs) with three-level CPU caching.
8. Limitations
Sisyphus is designed to optimize affine loop nests. For applications with non-affine code, we can partition the code into affine and non-affine sections. Sisyphus handles affine regions effectively, while tools like AutoDSE (Sohrabizadeh et al., 2021) or HARP (Sohrabizadeh et al., 2023) can optimize non-affine parts. The optimized sections are then integrated using a dataflow model, either with FIFOs for data exchange or by employing TAPA (Guo et al., 2023), which facilitates dataflow programming. This combined approach ensures that both affine and non-affine portions of the application are handled in an optimal manner, achieving a balance between different types of optimizations and seamlessly integrating them for improved overall performance.
@@ 9. Conclusion
In this paper, we introduced Sisyphus, a unified framework that streamlines hardware design through high-level synthesis. By integrating code transformation, pragma insertion, and tile size selection into a single optimization, Sisyphus addresses challenges in existing HLS workflows, which often rely on manual and disjointed methods. Sisyphus defines a design space in the form of a template. Using NLP, it efficiently explores this space, quickly identifying transformations and pragma configurations to achieve designs with good QoR. Our evaluation demonstrates that Sisyphus outperforms AutoDSE, NLP-DSE, and ScaleHLS in terms of QoR.